The demand for product X depends on the price of product X as well as the average household income (Y) according to the following relationship

Qdx = 600 - 25 P + 0.001Y


The supply of product X is positively related to own price of product X and negatively dependent upon W, the price of some input. This relationship is expressed as:

Qsx = 130 + 20 P - 3 W

Given that Y = 25,000 and W = 7, what is the:

1. Equilibrium price?    

2. Equilibrium quantity?    

Suppose that income increases to 35,000 and W remains constant. What is the new:

3. Equilibrium price?    

4. Equilibrium quantity?    

Assuming that income remains constant at 35,000 and W increases to 12, what is the new:


5: Equilibrium price?    

6. Equilibrium quantity?    
Answer #1

Qdx = 6oo-250 +0.ooly Qsu = 130+ 2op 3w. ( 4 = 25000 W=7. 6 dans 600-25P+0.001(25000) 600-25P +25 TQdx = 625-250 Q sa = 130.7